OpenStudy (anonymous):

Solve Please: 6x=9+x^2

OpenStudy (anonymous):

you just make it an equation and solve...so itz (x-3)^2 lol

OpenStudy (diyadiya):

x^2-6x+9=0 x^2-3x-3x+9=0 x(x-3)-3(x-3)=0 (x-3)(x-3)=0 x-3=0 x=3 :)

OpenStudy (anonymous):

Thank you very much!:D

OpenStudy (diyadiya):

:D